# When to Localize? A Risk-Constrained Reinforcement Learning Approach

### v1

@@ Abstract
In a standard navigation pipeline, a robot localizes at every time step to lower navigational errors. However, in some scenarios, a robot needs to selectively localize when it is expensive to obtain observations. For example, an underwater robot surfacing to localize too often hinders it from searching for critical items underwater, such as black boxes from crashed aircraft. On the other hand, if the robot never localizes, poor state estimates cause failure to find the items due to inadvertently leaving the search area or entering hazardous, restricted areas. Motivated by these scenarios, we investigate approaches to help a robot determine “when to localize?” We formulate this as a bi-criteria optimization problem: minimize the number of localization actions while ensuring the probability of failure (due to collision or not reaching a desired goal) remains bounded. In recent work, we showed how to formulate this active localization problem as a constrained Partially Observable Markov Decision Process (POMDP), which was solved using an online POMDP solver. However, this approach is too slow and requires full knowledge of the robot transition and observation models. In this paper, we present RiskRL, a constrained Reinforcement Learning (RL) framework that overcomes these limitations. RiskRL uses particle filtering and recurrent Soft Actor-Critic network to learn a policy that minimizes the number of localizations while ensuring the probability of failure constraint is met. Our numerical experiments show that RiskRL learns a robust policy that outperforms the baseline by at least 13% while also generalizing to unseen environments.
@@ I Introduction
In robotics, self-localization is crucial because it enhances navigation accuracy, and situational awareness, and enables complex tasks. Typically, an autonomous robot perceives its environment and self-localizes, plans its subsequent actions, acts upon its plan phases, and repeats the cycle. However, sometimes, a robot may want to localize seldom when it is not advantageous. For example, underwater robots need to surface to localize in underwater rescue and recovery missions. Surfacing to localize too often may hinder an underwater robot from searching for critical underwater items such as black boxes from crashed aircraft. On the other hand, if the robot never localizes, it will accumulate large amounts of drift [1], which may prevent it from finding the items due to inadvertently leaving the search area or entering hazardous, restricted areas. Therefore, in such scenarios, robots must balance prolonged actions that achieve mission objectives (such as searching for critical items underwater) with localizing to improve navigation accuracy.
@@
Although our prior approach produced policies that reached the goal and outperformed baselines, the approach had limitations. First, CC-POMCP was computationally expensive, requiring over 20 minutes of inference to navigate a path of 55 waypoints [^1] We obtained these run times from performing the experiments on a computing cluster.. Second, CC-POMCP requires a well-defined model of the environment, including the robot’s transition (motion) and observation models. Requiring such models may be problematic in unknown or dynamic real-world environments where it may be challenging to obtain accurate models. Finally, in our prior experiments, CC-POMCP failed to reach the goal when we increased the amount of transition noise, which limited successful experiments to low transition noises.
We propose a novel approach termed RiskRL that employs constrained Reinforcement Learning (RL) [4] and Particle Filters (PF) to overcome these limitations. Our new approach has multiple advantages over our prior work. First, although our new approach has a longer, single training time, it infers quicker during deployment, enabling real-time planning. The second was reducing the need for accurate transition and observation models of the environment. Like cost-constrained POMDPs in our prior work [2], constrained RL allows us to separate the value function and risk. In this formulation, the risk is modeled as a probability constraint, allowing us to design policies that minimize the failure probability while ensuring the robot remains within acceptable risk levels. This formulation provides greater control over the failure rate by explicitly incorporating risk constraints into the decision-making process. Furthermore, we use a PF to maintain the robot’s belief as the robot executes noisy motion commands and receives noisy measurements from the environment. We also use the PF to compute the observation for the RL robot.
We perform numerical experiments to compare RiskRL with several baselines, including standard RL (BaseRL), CC-POMCP, and heuristic policies. Our main finding is that when deployed in the environment that the robot is trained in RiskRL outperforms the BaseRL and CC-POMCP baselines by at least 13% in terms of the success rate while also being the only algorithm that satisfies the risk constraint. While environment-specific heuristics can match the performance, we find that RiskRL also generalizes to unseen environments, which the heuristics cannot.
@@ II-A Active Localization
Active perception [5, 6] encompasses various approaches, including active localization [7, 8], active mapping [9, 10, 11], and active SLAM [9]. These approaches focus on finding optimal robot trajectories and observations to achieve mission goals. Of these approaches, our current approach falls under active localization. Typically, active localization methods address where a robot moves and looks to localize itself [12, 13, 14, 15] or a target [16, 17]. Thus, these active localization approaches generally differ from our problem because we seek to determine when a robot localizes. However, one exception is our prior work [2], which proposes a preceding approach to the one in this paper. This approach improves upon [2], where we now model the probability of failure in terms of risk, reduce the inference time significantly, and relax the need for well-defined noise models.
@@ II-B Constrained Reinforcement Learning
Figure 2: This flowchart depicts a decision-making process for a robot interacting with an environment. It integrates a particle filter for state estimation and reinforcement learning components (actor-critic architecture) for decision-making.
@@ IV RiskRL Active Localization Algorithm
In this section, we present our RiskRL algorithm for solving the active localization problem given in Equation 1. Within the context of RiskRL, the following subsections discuss our active localization framework (that is, the PF, low-level planner, and high-level RL planner) in Section IV-A and provide detailed descriptions of the RL algorithm in the remaining subsections.
IV-A Overview
@@
We now discuss our RiskRL active localization framework, which is depicted in Figure 2 and described in Algorithm 1. Our framework comprises three main components: an RL high-level planner, a PF, and a low-level planner. The high-level planner chooses which action (move or localize) the agent performs given an input state $s_{t}$. Next, we use the PF to maintain a belief of the agent’s state whenever the agent selects either action and computes a state vector for the RL planner. Finally, a low-level planner determines how the agent should move (that is, move forward, turn left, or turn right) if move is chosen or recomputes hazard-free paths if localize is chosen.
IV-B High-Level Policy
The output from the PF is fed as input to the high-level policy. The goal of the high-level policy is to choose between two actions: $\mathbb{A}_{\text{RLP}}=\{\text{move},\text{localize}\}$. If the policy chooses to move, then we execute the next action given by the low-level planner without taking an observation. As described earlier, we use a constrained RL policy that minimizes the number of localization actions while ensuring the risk constraint is met.
The state input $s_{t}$ given to the RL policy is derived from the output of the PF. Specifically, $s_{t}$ is a 2D vector containing the collision probability and the Manhattan distance from the robot to the goal. We compute the collision probability by counting the number of particles in collision at the current time step. We chose this observation representation so that our agent generalizes to unseen environments.
The reward $r_{t}=-1$ if the robot chooses the localize action and 0 otherwise. We can use standard RL to minimize the number of localization actions by maximizing this reward function. However, naive optimization will violate the constraints in (1). The constraint probability is difficult to estimate because it requires interaction with the environment and varies based on the policy being used, making it challenging to establish a clear relationship between the policy and the constraint. We follow the relaxation approach outlined in [4], converting the probabilistic constraint in a Chance-Constrained POMDP into a cumulative constraint. Specifically, we reformulate the optimization problem as follows. Our goal is to maximize the expected cumulative reward $V(\theta)$, defined by:
@@ IV-B High-Level Policy
where $\theta$ denotes the parameters of the policy $\pi_{\theta}$, and $V(\theta)$ represents the expected reward over time. We then impose a cumulative constraint:
$U_{\theta}:\sum_{t=0}^{T}\gamma^{t}(1-\Pr(\text{failure}\mid x_{1:N},a_{1:t},b(s_{0})))\geq c$ (3)
Here, $U_{\theta}$ is the accumulated discounted probability of failure, and $c=\frac{(1-\hat{c}\gamma^{T}(1-\gamma))}{(1-\gamma)}$ represents the risk-adjusted threshold. This formulation simplifies the original problem by approximating the probabilistic constraint, thus allowing the constraint to be evaluated based on the data generated from the rollout trajectory.
@@
The algorithm [4] shown below employs the primal-dual method to optimize the expected reward while satisfying constraints. The primal component focuses on maximizing the reward function as defined in Equation (4), while the dual component adjusts the $\lambda$ values to control the risk levels associated with these constraints. It iteratively updates the policy parameters and dual variables by simulating trajectories and estimating gradients. The policy can be updated by computing the policy gradient defined by the following equation:
$\nabla_{\theta}L(\theta_{k},\lambda_{k})=\hat{r}(s_{t},a_{t})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{0}|s_{0})$ (5)
where $\nabla_{\theta}\log\pi_{\theta_{k}}(a_{0}|s_{0})$ is the gradient of the log-probability of the policy $\pi_{\theta_{k}}$ selecting action $a_{0}$ given state $s_{0}$.
@@
IV-C Handling Partial Observability
The previous section described how we can optimize the policy using the primal-dual approach. This section presents the specific architecture we use for the actual policy. Stemming from [21], we use a Soft Actor-Critic Model as it generally tends to have better sample efficiency. To deal with partial observability, Recurrent Neural Networks (RNN) have been known to mitigate the effects of a noisy observation by making decisions based on the past trajectory instead of just the current observation (or fixed sequence of observations) [25][26]. We implement an LSTM module to help stabilize training [27]. Initial training results showed that a deterministic actor would more often converge to a uniform policy (only moving or only localizing), therefore we decided to optimize the results with a stochastic actor. All of our embeddings are obtained with a one-layer MLP. Figure 3 provides an illustrative diagram of the architecture.
@@ V Experimental Evaluation
In this section, we report our findings from numerical experiments comparing RiskRL with several baselines and evaluating the robustness and generalization capabilities of RiskRL. The training and testing environment are motivated by an underwater scenario introduced in our prior work [2]. In the scenarios, a robot is tasked with navigating through complex underwater environments, which include obstacles such as rocks and coral formations. The robot must perform localization and path planning while accounting for the unique challenges posed by water currents and variable visibility. Our experimental code is located on GitHub [^2] https://github.com/raaslab/when-to-localize-riskrl.
@@ V-A Setup
Figure 4: This figure shows the three Minigrid environments used in our evaluations. The red triangles represent the robots, the green squares represent the goal, and the orange squares represent the failure states. We trained the RL robots using the train environment. All environments are in a 32x32 grid.
Baselines. We compare RiskRL with four types of baselines. The first type is two static policies (SP): (M2x,L), and (M3x,L), where for example, (M2x,L) is a fixed repeating sequence of two move actions followed by a localize action. The second type is a threshold planner (TP) that localizes the robot whenever the probability of failure (computed using the PF output) exceeds a threshold. The third type is an implementation of CC-POMCP, a cost-aware, online policy planner from our prior work [2]. Finally, the last type is a standard, risk-unaware RL policy (BaseRL). In the latter, we penalize every time the robot chooses a localize action ($-1$) and when it collides ($-256$). This baseline is used to assess the advantage of employing risk-aware RL.
@@
TABLE I: RL Parameters. $\ell_{r}$ is the learning rate, $\gamma$ is the discount factor, $\alpha$ is the target entropy, and $\tau$ is the soft update parameter
Environments. Figure 4 shows our environmental setup. We assume the robot knows the map, the start state, and the goal region. We set the initial belief to the start state. In all of our experiments, we set the transition noise such that the robot has an 80% chance of going in the direction it is facing, a 10% chance of drifting to the left, and a 10% chance of drifting to the right. Other than the results in Section V-E, we assume no observation noise when localizing to focus more on the effect of transition noise. Figure 4 visually represents our environments.
Training Process. We trained each of the planners on the train environment, allotting 8 hours for BaseRL and 12 hours for RiskRL to reach  200 episodes. The parameters of the BaseRL robot and the proposed RiskRL robot are defined in Table I.
Figure 5: A qualitative example of robot navigation: blue) true path, red) estimated path and localization points.
The following subsections compare the performance of our baselines and the proposed algorithm. Our experiments ran until the robot reached the goal or failure region. The results were based on 100 runs for each algorithm in each environment (except CC-POMCP, which we only ran 75 times in the train environment since it takes around 30X more time to get through a trajectory compared to the rest of the planners). All bars depict the mean values in Figures 7(a) and 7(b).
V-B Qualitative Example
In Figure 5, we present a qualitative example demonstrating the robot’s behavior in a noisy and uncertain environment, characterized by (80%, 10%, 10%) transition noise and 32% observation noise, with a risk threshold set at 40%. The results show that the robot consistently applies localization in the Start Area and Middle Tunnel to mitigate the risk of early failure. Additionally, the robot increases the frequency of localization within the Middle Tunnel, where the risk of failure is higher, and noise can cause deviations from the intended path. Since the area after the Middle Tunnel is a safe zone, the robot does not perform localization there. Finally, the robot localizes upon reaching the goal to ensure precise positioning for successful task completion. This adaptive behavior highlights the robot’s strategy to navigate effectively under challenging conditions.

### v2

@@ Abstract
In a standard navigation pipeline, a robot localizes at every time step to lower navigational errors. However, in some scenarios, a robot needs to selectively localize when it is expensive to obtain observations. For example, an underwater robot surfacing to localize too often hinders it from searching for critical items underwater, such as black boxes from crashed aircraft. On the other hand, if the robot never localizes, poor state estimates cause failure to find the items due to inadvertently leaving the search area or entering hazardous, restricted areas. Motivated by these scenarios, we investigate approaches to help a robot determine “when to localize?” We formulate this as a bi-criteria optimization problem: minimize the number of localization actions while ensuring the probability of failure (due to collision or not reaching a desired goal) remains bounded. In recent work, we showed how to formulate this active localization problem as a constrained Partially Observable Markov Decision Process (POMDP), which was solved using an online POMDP solver. However, this approach is too slow and requires full knowledge of the robot transition and observation models. In this paper, we present RiskRL, a constrained Reinforcement Learning (RL) framework that overcomes these limitations. RiskRL uses particle filtering and recurrent Soft Actor-Critic network to learn a policy that minimizes the number of localizations while ensuring the probability of failure constraint is met. Our numerical experiments show that RiskRL learns a robust policy that leads to at least a 26% increase in success rates when traversing unseen test environments.
Code: https://github.com/raaslab/when-to-localize-riskrl
@@ I Introduction
In robotics, self-localization is crucial because it enhances navigation accuracy, and situational awareness and enables complex tasks. Typically, an autonomous robot perceives its environment and self-localizes, plans its subsequent actions, acts upon its plan phases, and repeats the cycle. However, sometimes, a robot may want to localize seldom when it is not advantageous. For example, underwater robots need to surface to localize in underwater rescue and recovery missions. Surfacing to localize too often may hinder an underwater robot from searching for critical underwater items such as black boxes from crashed aircraft. On the other hand, if the robot never localizes, it will accumulate large amounts of drift [1], which may prevent it from finding the items due to inadvertently leaving the search area or entering hazardous, restricted areas, as illustrated in Figure 1. Additionally, since the robot cannot execute movement and localization actions simultaneously, robots must balance prolonged actions that achieve mission objectives (such as searching for critical items underwater) with localizing to improve navigation accuracy.
@@
Although our prior approach produced policies that reached the goal and outperformed baselines, the approach had limitations. First, CC-POMCP was computationally expensive, requiring over 20 minutes of inference to navigate a path of 55 waypoints. Second, CC-POMCP requires a well-defined model of the environment, including the robot’s transition (motion) and observation models. Requiring such models may be problematic in unknown or dynamic real-world environments where it may be challenging to obtain accurate models. Finally, CC-POMCP is highly sensitive to transition noise, often failing to reach the goal as noise increases. As shown in Figure 6, RiskRL achieves a higher success rate under the same transition noise conditions, demonstrating greater robustness.
We propose a novel approach termed RiskRL that employs constrained Reinforcement Learning (RL) [4] and Particle Filters (PF) to overcome these limitations. Our new approach has multiple advantages over our prior work. First, although our new approach has a longer, single training time, it infers quicker during deployment, enabling real-time planning. The second was reducing the need for accurate transition and observation models of the environment, which we demonstrate by varying the transition and observation noises. In the formulation, the risk is modeled as a probability constraint, allowing us to design policies that minimize the failure probability while ensuring the robot remains within acceptable risk levels. This formulation provides greater control over the failure rate by explicitly incorporating risk constraints into the decision-making process. Furthermore, we use a PF to maintain the robot’s belief as the robot executes noisy motion commands and receives noisy measurements from the environment. We also use the PF to compute the observation for the RL robot.
We perform numerical experiments to compare RiskRL with several baselines, including standard RL (BaseRL), CC-POMCP, and heuristic policies. Our main finding is that when deployed in unseen testing environments, RiskRL outperforms the BaseRL and CC-POMCP baselines by at least 26% in terms of the success rate while also being the only algorithm that satisfies the risk constraint. Unlike BaseRL, CC-POMCP, and the heuristics algorithm, RiskRL generalizes to our unseen test environments.
The remainder of this paper is organized as follows: Section II reviews the related work in active localization and constrained RL. Section III defines the problem formulation. Section IV presents the proposed RiskRL framework. Section V provides experimental evaluations and comparisons with baselines. Finally, Section VI concludes the paper and discusses future directions.
@@ II-A Active Localization
Active perception [5, 6] encompasses various approaches, including particle filter [30, 29], active localization [7, 8], active mapping [9, 10, 11], and active SLAM [9]. These approaches focus on finding optimal robot trajectories and observations to achieve mission goals. Of these approaches, our current approach falls under active localization. Typically, active localization methods address where a robot moves and looks to localize itself [12, 13, 14, 15] or a target [16, 17]. Thus, these active localization approaches generally differ from our problem because we seek to determine when a robot localizes. However, one exception is our prior work [2], which proposes an approach that precedes the one in this paper. This approach improves upon [2], where we now model the probability of failure in terms of risk, reduce the inference time significantly, and relax the need for well-defined noise models.
@@ II-B Constrained Reinforcement Learning
Figure 2: This flowchart depicts a decision-making process for a robot interacting with an environment. It combines a high-level (Soft Actor-Critic) RL planner for decision-making (Section IV-A) and a Particle Filter for state estimation.
@@ IV RiskRL Active Localization Algorithm
We now present our algorithmic framework (Figure 2) for solving the active localization problem in (1). Our framework has three main components: a high-level planner, a low-level planner, and a PF. We discuss the high-level planner (which chooses when to move or localize) and our RL solution in Section IV-A. The low-level planner selects a motion command $u_{t}$ if the agent wants to move, or localizes and replans its trajectory if it wants to localize. Finally, the PF provides observations $o_{planner}$ to the high-level planner and maintains the agent’s belief using $u_{t}$ and noisy, 2D pose observations $o_{state}$ from the environment. Algorithm 1 describes how our active localization algorithm works.
IV-A High-Level Planning
Overview. The high-level planner chooses which high-level action the agent performs next: $a\in\mathbb{A}_{\text{RLP}}=\{\text{move},\text{localize}\}$. The planner chooses the next action based on the 2D observation vector $o_{planner}$ from the PF. This vector contains the predicted collision probability $\hat{p}\in[0,1]$ and distance $\hat{d}\in[0,+\inf)$ to the goal. We compute $\hat{p}$ by counting the number of particles that collided with an obstacle and $\hat{d}$ as the number of steps between the belief’s mean and the goal.
In this paper, we implement the high-level planner using heuristics for our baselines without learning and SAC neural networks for RiskRL and BaseRL. We describe the heuristics-based planners in Section V and the RiskRL planner below. To select the next high-level action $a_{t}$, the RL planners use the hidden state $h_{t}$ computed from the LSTM module, the previous action $a_{t-1}$, the observation vector $o_{planner}$, while the heuristic planners either use $o_{planner}$ alone or internal data such as the number of moves between localize actions. To train the RL agent, the reward $r_{t}$ is chosen to be $-1$ if the robot chooses the localize action and 0 otherwise. We chose this observation representation so that our RL planner generalizes to unseen environments.
Chance-Constrained Planning (RiskRL). The original objective (1) does not align with the standard RL formulation for expected discounted reward. Thus, we propose a relaxation, as shown in Equation 2. This relaxation allows us to use standard RL to minimize the number of localization actions by maximizing this reward function. However, naive optimization will violate the constraints in (1).
The constraint probability is difficult to estimate because it requires interaction with the environment and varies based on the policy being used, making it challenging to establish a clear relationship between the policy and the constraint. We follow the relaxation approach outlined in [4], converting the probabilistic constraint in a Chance-Constrained POMDP into a cumulative constraint. Specifically, we reformulate the optimization problem as follows. Our goal is to maximize the expected cumulative reward $V(\theta)$, defined by:
@@ IV-A High-Level Planning
where $\theta$ denotes the parameters of the policy $\pi_{\theta}$ and $V(\theta)$ denotes the expected reward over time. We then impose a cumulative constraint:
$U_{\theta}:\sum_{t=0}^{T}\gamma^{t}(1-\Pr(\text{failure}\mid x_{1:N},a_{1:t},b(s_{0})))\geq c,$ (3)
where $U_{\theta}$ is the accumulated discounted probability of failure and $c=\frac{(1-\hat{c}\gamma^{T}(1-\gamma))}{(1-\gamma)}$ represents the risk-adjusted threshold. This formulation simplifies the original problem by approximating the probabilistic constraint, allowing us to evaluate the constraint based on the data generated from the rollout trajectory.
@@
The algorithm [4] shown below employs the primal-dual method to optimize the expected reward while satisfying constraints. The primal component focuses on maximizing the reward function as defined in Equation (4), while the dual component adjusts the $\lambda$ values to control the risk levels associated with these constraints. It iteratively updates the policy parameters and dual variables by simulating trajectories and estimating gradients. The policy can be updated by computing the policy gradient defined by ${\nabla_{\theta}L(\theta_{k},\lambda_{k})=\hat{r}(s_{t},a_{t})\nabla_{\theta}\log\pi_{\theta_{k}}(a_{0}|s_{0})}$, where $\nabla_{\theta}\log\pi_{\theta_{k}}(a_{0}|s_{0})$ is the gradient of the log-probability of the policy $\pi_{\theta_{k}}$ selecting action $a_{0}$ given state $s_{0}$.
@@
Handling Partial Observability. The previous section described how we can optimize the policy using the primal-dual approach. This section presents the specific architecture we use for the actual policy. Stemming from [21], we use a Soft Actor-Critic Model as it generally tends to have better sample efficiency. To deal with partial observability, Recurrent Neural Networks (RNN) have been known to mitigate the effects of a noisy observation by making decisions based on the past trajectory instead of just the current observation (or fixed sequence of observations) [25][26]. We implement an LSTM module to help stabilize training [27]. All of our embeddings are obtained with a one-layer MLP. Figure 3 provides an illustrative diagram of the architecture.
@@ V Experimental Evaluation
In this section, we report our findings from numerical experiments comparing RiskRL with several baselines and evaluating the robustness and generalization capabilities of RiskRL. The training and testing environment are motivated by an underwater scenario introduced in our prior work [2]. In such work, a robot must balance localizing at the surface and navigating through underwater environments, which include obstacles such as rocks and coral formations, to search for items such as the black boxes of downed airplanes.
@@ V-A Setup
Figure 4: This figure shows our three evaluation environments in Minigrid. A red triangle denotes the agent, the orange squares denote the failure states (obstacles), and the sample paths are denoted in blue with the green squares representing the goal. We trained the RL agents using train.
Baselines. We compare RiskRL with four types of baselines. The first type is two static policies (SP): (M2x,L), and (M3x,L), where (M2x,L) repeatedly moves twice and then localizes once. The second type is a threshold planner (TP) that localizes the robot whenever the collision probability exceeds a threshold. The third type is CC-POMCP, a cost-aware, online policy planner from our prior work [2]. Finally, the last type is BaseRL, a standard, risk-unaware RL policy. For BaseRL, we penalize the robot each time it localizes or collides (Table I). This baseline is used to assess the advantage of employing risk-aware RL.
@@
TABLE I: RL Parameters. $\ell_{r}$ is the learning rate, $\gamma$ is the discount factor, $\alpha$ is the target entropy, and $\tau$ is the soft update.
Environments. Figure 4 shows our environmental setup. We assume the robot knows the map, the start state, and the goal region. We set the initial belief to the start state. In all of our experiments, we set the transition noise such that the robot has an 80% chance of going forward and a 10% chance of drifting to the left or the right. The observation noise is drawn from a 3x3 matrix with a 68% probability of observing its true position and a 4% probability of observing a neighboring position. We assume no observation noise when localizing to focus more on the effect of transition noise, except for the results in Section V-E.
Training Process. We trained each planner in the train environment, allotting 8 hours for BaseRL and 12 hours for RiskRL to reach  200 episodes. Each model was trained using a GTX 2080 Ti. The parameters of the BaseRL and RiskRL agents are defined in Table I.
Figure 5: A qualitative example of robot navigation: blue) actual path, red) estimated path and localization positions.
The following subsections compare the performance of our baselines and RiskRL. Our experiments ran until the robot reached the goal or failure region. The results were based on 100 runs for each algorithm in each environment (except CC-POMCP, which only has 75 runs in the train environment). In the chesapeake and st_thomas environments, we randomized the path for each iteration, where the start and goal points were randomly selected from two open spaces that were at least 50 blocks away from each other.
V-B RiskRL Qualitative Example
Figure 5 presents a qualitative example demonstrating the robot’s behavior in a noisy and uncertain environment, characterized by a (80%, 10%, 10%) transition noise, (68%, 4%) observation noise, and 40% risk threshold. The results show that the robot consistently localizes in the Start Area and Middle Tunnel to mitigate failing early. Additionally, the agent increases its localization frequency within the Middle Tunnel, where the risk of failure is higher and noise can cause deviations from the intended path. Since the area after the middle tunnel is more spacious and thus safer, the agent does not localize as often. Finally, the agent localizes near the goal to ensure precise positioning for successful task completion. This adaptive behavior highlights the agent’s strategy to navigate effectively under challenging conditions.